OpenStudy (anonymous):

The vertices of triangle ABC are A(0,6) B(3,0) C(11,0) What is the area of triangle ABC in square units? A) 24 B) 33 C) 12 D) 9

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

area of triangle = (1/2) base*height base = 11-3 = 8 height = 6-0 = 6

OpenStudy (anonymous):

48/2=24

OpenStudy (anonymous):

A) 24

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thx